A dairy man wants to mix 85% protien supplement and a standard 15% protein ration to make 1700 pounds of higgh-grade 50% protein ration. How many pounds of each shoul he use?
Answer by checkley75(3666)

.85X+.15(1700-X)=.5*1700
.85X+255-.15X=850
.7X=850-255
.7X=595
X=595/.7
X=850 POUNDS OF 85% PROTIEN IS NEEDED.
1700-850=850 POUNDS OF 15% PROTIEN IS NEEDED.
PROOF:
.85*850+.15*850=.5*1700
722.5+127.5=850
850=850